OpenStudy (anonymous):

can anyone help me understand this problem Solve using the addition principle. Graph and write set-builder notation and interval notation for each answer. 5(t + 3) + 9 ≥ 3(t − 2) − 10 i have solved it but not sure if i got it right {t|t ≥ 0} or (–∞, 0]

OpenStudy (anonymous):

what class is it for. Algebra or set theory?

OpenStudy (anonymous):

basic algebra

OpenStudy (anonymous):

Can you tell me how you arrived at that solution?

OpenStudy (anonymous):

5(t + 3) + 9 ≥ 3(t − 2) − 10 5t+3+9 ≥ 3t-2-10 5t+12 ≥ 3t-12 5t-3t +12-12 ≥ 3t-3t-12+12 2t /2 ≥ 0/2 t≥0 that was the answer i got

OpenStudy (anonymous):

I think I see what you did wrong

OpenStudy (anonymous):

5(t+3)+9>or= 3(t-2)-10 You have to use distributive property.(Everything inside the parenthesis should be multiplied by number outside.)

OpenStudy (anonymous):

dam it cant believe i let that one by me

OpenStudy (anonymous):

So it is 5*t+5*3+9>0r=3*t+(3*-2) -10

OpenStudy (anonymous):

So did you end up getting an answer?

OpenStudy (anonymous):

working it out now

OpenStudy (anonymous):

5(t + 3) + 9 ≥ 3(t − 2) − 10 5t+15+9 ≥ 3t-6-10 5t+24 ≥ 3t-16 5t-3t +24-24 ≥ 3t-3t-16+24 2t /2 ≥ 8/2 t≥4 {t|t ≥ 4} or (–∞, 4]

OpenStudy (anonymous):

not sure about the last line though thats whats messing me up

OpenStudy (anonymous):

since t is greater or equal to 4 we should start from 4 to infinity. so it should be [4,infinity)

OpenStudy (anonymous):

ok so i have the answer backwards 5(t + 3) + 9 ≥ 3(t − 2) − 10 5t+15+9 ≥ 3t-6-10 5t+24 ≥ 3t-16 5t-3t +24-24 ≥ 3t-3t-16+24 2t /2 ≥ 8/2 t≥4 {t|t ≥ 4} or [4,∞)

OpenStudy (anonymous):

thank you i see im going to hate my next math class more then this one